OpenStudy (anonymous):

what is equivalent to the following equation y=x^2+4x+11

OpenStudy (amistre64):

the compressed form of it i would assume

OpenStudy (amistre64):

add 4 and subtract 4 to compress it

OpenStudy (anonymous):

add 4 where and subtract where?

OpenStudy (amistre64):

y=x^2+4x+(4-4)+11 y=(x^2+4x+4)-4+11 y=(x+2)^2 +7

OpenStudy (amistre64):

in other words, the method is generally refered to as "completeing the square" if you wanna read up on it

OpenStudy (anonymous):

So y=(x+3)^2+53 would be the same way?

OpenStudy (amistre64):

if they want a different form, then that one would need to be expanded

OpenStudy (amistre64):

(x+2) * (x+3) = x^2 +6x + 9 y = x^2 +6x + 9 + 53

OpenStudy (anonymous):

Thanks